a violinist plays a note of 440 Hz on his violin. How long does it take the violin string to move one complete back and forth cycle? (enter answer in seconds)
Answer #1

Solution:

Given:

Frequency (f) = 440 Hz

We need to find Period of the cycle:

Since: Period (T) = 1 / f = 1 / (440 Hz) = 0.00227 sec

Therefore: time taken by the violin string to move one complete back and forth cycle is 0.00227 sec.